what are the dimensions of a rectangle that has a perimeter of 74 ft and an area of 36ft^2?
Answer by checkley77(12844)

XY=36 OE X=36/Y
2X+2Y=74
2(36/Y)+2Y=74
72/Y+2Y=74
(72+2Y*Y)/Y=74
(72+2Y^2)/Y=74 CROSS MULTIPLY
72+2Y^2=74Y
2Y^2-74Y+72=0
2(Y^2-37Y+36)=0
2(Y-36)(Y-1)=0
Y-36=0
Y=36 ANS.
X=36/36=1 ANS.
-------------
Y-1=0
Y=1 ANS
X=36/1=36 ANS
PROOFS:
1*36=36=36
2*36+2*1=74
2+72=74
74=74